OpenStudy (anonymous):

how to I simplify p^(-2/3)/ p^(5/6)?

OpenStudy (anonymous):

I have the answer for this one -> 1/p^(1/6) but I don't know how to get to that answer...

OpenStudy (anonymous):

Subtract the second exponent from the first.

OpenStudy (anonymous):

Subtract the exponents.

OpenStudy (anonymous):

that doesn't get me to the right answer...

OpenStudy (anonymous):

\[p^{-\frac{2}{3}} / p^{\frac{5}{6}} = p^{-\frac{2}{3} - \frac{5}{6}}\]\[=p^{-\frac{3}{2}} = 1/{p^{\frac{3}{2}}}\] That's what I get.

OpenStudy (anonymous):

okay. that's not the right answer but it helped me get it. thanks

OpenStudy (anonymous):

Coincidentally, this gives the answer you have.. \[p^\frac{2}{3}/ p^\frac{5}{6}\]

OpenStudy (anonymous):

The answer I gave is definitely correct.